OpenStudy (deaaan):

Simplify (x − 3)(x2 + 7x − 8).

OpenStudy (madisongoode):

So: 1.1 Factoring x2+7x-8 The first term is, x2 its coefficient is 1 . The middle term is, +7x its coefficient is 7 . The last term, "the constant", is -8 Step-1 : Multiply the coefficient of the first term by the constant 1 • -8 = -8 Step-2 : Find two factors of -8 whose sum equals the coefficient of the middle term, which is 7 . -8 + 1 = -7 -4 + 2 = -2 -2 + 4 = 2 -1 + 8 = 7 That's it Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, -1 and 8 x2 - 1x + 8x - 8 Step-4 : Add up the first 2 terms, pulling out like factors : x • (x-1) Add up the last 2 terms, pulling out common factors : 8 • (x-1) Step-5 : Add up the four terms of step 4 : (x+8) • (x-1) Which is the desired factorization Final result : (x - 3) • (x + 8) • (x - 1) Does this help?